OpenStudy (anonymous):

What is the y-intercept of the equation? y = 5(6.8)^x A. 5 B. 6.8 C. 1

Directrix (directrix):

A)5

Directrix (directrix):

y = 5(6.8)^x The y-intercept is found by setting x = 0. y = 5* (6.8) ^ 0 y = 5 * 1 = 5